solve the following using a system of equations;
Rod has 40 coins, all dimes and quarters, worth $7.60. How many dimes and how many quarters does he have?

Answer by stanbon(75887)

Rod has 40 coins, all dimes and quarters, worth $7.60. How many dimes and how many quarters does he have?
--------
Equations:
Quantity Eq. d + q = 40
Value Eq:: 10d+25q = 760
------------------------------
Multiply thru the Quantity equation by 10 to get:
10d + 10q = 10*40
-----
Subtract that from the Value Eq and solve for "q":
15q = 360
q = 24 (# of quarters)
-----
Substitute into d+q = 40 and solve for "d":
d + 24 = 40
d = 16 (# of dimes)
=========================
Cheers,
Stan H.